OpenStudy (anonymous):

Simplify the expression. -l-13l = ???

OpenStudy (anonymous):

You are being asked what is the absolute value of -13 and then to negate that value.

OpenStudy (ghazi):

I-aI=a

OpenStudy (anonymous):

y questions says SIMPLFIY

OpenStudy (anonymous):

my

OpenStudy (anonymous):

Because you can simplify it to -13.

OpenStudy (ghazi):

after that you have to find the final value of it

OpenStudy (anonymous):

omg

OpenStudy (ghazi):

I-aI=a but -I-aI=-a

OpenStudy (anonymous):

understand? I-13I=13, but -I-13I=-13